a university as 2000 male and 500 female faculty members. a sample of 50 female and 200 male faculty members gives each member of the faculty member 1 chance in 10 to be chose. this sample design gives every unit in the population the same chance to be chosen for the sample. is it a simple random sample? explain your answer.
1
Expert's answer
2012-02-03T08:23:40-0500
Yes it is simple random sample- quota sample, because chanses to be choosen :
Women 50/500=0.1
Men& 200/2000=0.1
P(to choose men)=2000/2500=4/5
P(to choose& women)= 500/2500=1/5
Chanses to be choosen by formula of full probability
P(do choose a different person)=P(to be choosen among men)*P(to choose men)+ P(to be choosen among women)*P(to choose women)= 0.1*4/5+0.1*1/5=0.1=
= P(to choose men)= P(to choose& women).